OpenStudy (anonymous):

Give an example of a rational function that has a horizontal asymptote at y = 1 and a vertical asymptote at x = 4.

OpenStudy (anonymous):

okay you just have to find a function who tends to infinity at x=4 and when x tends infinity y=1 that's it

OpenStudy (anonymous):

how do you do that?

OpenStudy (anonymous):

okay let me have a equation y=tanx as you go to x=pi/2 the y approaches to infinity so we say at x=pi/2 we have a vertical asymptote .. are you getting what i'm saying?

OpenStudy (anonymous):

yes kind of.

OpenStudy (anonymous):

wait a sec please

OpenStudy (anonymous):

okay try this function 1/y-1=log(x-4) i think this should be answer to your question.. :)